A cyclist traveled at a rate of 24 mph to visit a nearby town. The cyclist averaged 18 mph on the return trip. If the round trip took 4.9 hours, find the distance to the nearby town? [I need the correct college algebra equation of the form D = rt]
Answer by checkley71(8403)

D=RT
SEEING AS THE DISTANCES ARE THE SAME WE HAVE:
24X=18(4.9-X)
24X=88.2-18X
24X+18X=88.2
42X=88.2
X=88.2/42
X=2.1 HOURS FOR THE 24 MPH TRIP.
4.9-2.1=2.8 HOURS FOR THE 18 MPH TRIP.
PROOF
24*2.1=18*2.8
50.4=50.4 THEREFORE THE DISTANCE TO THE NEARBY TOWN IS 50.4 MILES.